A body of mass 80kg moving with a velocity of 3 meter per second hit a stationary body of mass 40kg.

a. If the bodies move together after impact . Calculate the velocity of the composite body just after impact.

b. If the bodies do not move together and the body of mass 40kg is moving with a velocity of 3 meter per second in the same direction as that of the body of mass 80kg . Find the velocity of the 80kg after impact.


1
Expert's answer
2022-03-30T03:31:37-0400

a.

"u=v_1\frac{m_1}{m_1+m_2}=2~\frac ms."

b.

"u=v_1\frac{m_1-m_2}{m_1}=1.5~\frac ms."